A daredevil decides to jump a canyon. It's walls are equally high and 13 m apart. He takes off by driving a motorcycle up a short ramp sloped at an angle of 18°. What minimum speed must he have in order to clear the canyon?
Answer #1
Use range equation:
d = [v^2(sin2theta)]÷g
18 = [v^2(0.58778)]÷9.8
(18 x 9.8)÷0.58778 = v^2
v^2 = 216.745
v = 14.722 m/s

Hope this helps :)
answered by: Mira